OpenStudy (anonymous):

A man burns 10 calories per minute mountain climbing biking and 7.5 calories per minute in-line skating.His goal is to burn approximately 420 calories daily. Find the intercepts of the graph of the equation

OpenStudy (anonymous):

Anyone have any ideas?

OpenStudy (anonymous):

Are you having issues finding the equations to graph or the graph itself?

OpenStudy (anonymous):

both

OpenStudy (anonymous):

Okie dokie, so we have to find the equations first. He burns 10x calories for mountain biking (let x = number of minutes) and 7.5x for inline

OpenStudy (anonymous):

okay im listening...or readingg...

OpenStudy (anonymous):

Hmmm... it doesn't seem like they are going to intercept. Sorry, hang on. It's been awhile since I've done a problem like this. I'm going to find my old notes

OpenStudy (anonymous):

okay thanks i'll wait xD

OpenStudy (compassionate):

A man burns 10 calories per minute mountain climbing biking and 7.5 calories per minute in-line skating.His goal is to burn approximately 420 calories daily. Find the intercepts of the graph of the equation [(1, 10), (1, 7.5)] Plug those into the equation y - y1 = m(x - x1)

OpenStudy (compassionate):

Actually, first yo want to find the slope using the equation. m=(y2-y1)/(x2-x1) plug in the coordinates: [(1, 10), (1, 7.5)] Then plug [(1, 10), (1, 7.5)] and the newly found slope into: y - y1 = m(x - x1) Then put the equation in y = mx + b

OpenStudy (anonymous):

I don't think it can be solved that way... 10 and 7.5 aren't on the same line... Or am I looking at it wrong still?

OpenStudy (anonymous):

uuuuuhh! i honestly have no cllue

OpenStudy (anonymous):

Okay, I think Compassionate has the right idea, but 10 and 7.5 will the the slopes because that's how many calories they burn per minute.

OpenStudy (anonymous):

Are you sure there isn't more to the equation? because I'm only seeing them intersect at (0,0)

OpenStudy (anonymous):

well there's a question before it saying "Write an equation to represent the situation" so I am guessing we are suppose to make an equation first

OpenStudy (anonymous):

Okay, that's going to be y=10x and y=7.5x with y being the number of calories burned and x being the number of minutes

OpenStudy (compassionate):

No, the slope is found by using the two points above.

OpenStudy (compassionate):

m=(y2-y1)/(x2-x1) plug in the coordinates: [(1, 10), (1, 7.5)] to get your slope.

OpenStudy (anonymous):

But they aren't on the same line?

OpenStudy (anonymous):

For the slope I got 0

OpenStudy (anonymous):

He needs the intercept of the two lines and the values @Compassionate gave would give an undefined slope since you cannot divide by 0

OpenStudy (anonymous):

look this is the hw,its number 7

OpenStudy (anonymous):

@satellite73 can you help us with this? We are having trouble

OpenStudy (anonymous):

i think he/she is offline

OpenStudy (anonymous):

@CanadianAsian can you help us please?

OpenStudy (anonymous):

ugh, they were just online. @saifoo.khan can you maybe help us?

OpenStudy (anonymous):

Okay... Lets try to work this through step by step... did you already figure out which is the independent and which is the dependent variable?

OpenStudy (anonymous):

Yes,the independent variable are the minutes.The dependent variable are the calories

OpenStudy (anonymous):

Okay so x will be minutes and y will be calories (part of this I'm typing out for my own sake)

OpenStudy (anonymous):

part two makes it sound like there is only one equation they are looking for...

OpenStudy (compassionate):

... \[ \begin{array}l\color{red}{\text{(}}\color{red}{\text{U}}\color{red}{\text{S}}\color{red}{\text{E}}\color{red}{\text{R}}\color{red}{\text{ }}\color{red}{\text{W}}\color{red}{\text{A}}\color{red}{\text{S}}\color{red}{\text{ }}\color{red}{\text{B}}\color{red}{\text{A}}\color{red}{\text{N}}\color{red}{\text{N}}\color{red}{\text{E}}\color{red}{\text{D}}\color{red}{\text{ }}\color{red}{\text{F}}\color{red}{\text{O}}\color{red}{\text{R}}\color{red}{\text{ }}\color{red}{\text{T}}\color{red}{\text{H}}\color{red}{\text{I}}\color{red}{\text{S}}\color{red}{\text{ }}\color{red}{\text{P}}\color{red}{\text{O}}\color{red}{\text{S}}\color{red}{\text{T}}\color{red}{\text{)}}\color{red}{\text{}}\end{array} \]

OpenStudy (anonymous):

Yeah,that's what I got mostly confused on.I put that maybe it could be 10x+7.5y=420 idk

OpenStudy (anonymous):

Yeah, but that doesn't make sense though... That would be saying that 10* minutes + 7.5 * calories would equal 420 calories...

OpenStudy (anonymous):

Yeah that's true.

OpenStudy (anonymous):

Oh... they want the x and y intercepts of the equation. not where they intersect. That changes a lot. I'm a moron sometimes :P We still have to figure out the equation

OpenStudy (anonymous):

Lets try 10x +7.5x = y

OpenStudy (anonymous):

Haha no your not xD then we would be left with 17.5=7

OpenStudy (anonymous):

i mean 17.5=y

OpenStudy (anonymous):

17.5x=y yes

OpenStudy (anonymous):

17.5x=y there we go!

OpenStudy (anonymous):

I'm still not convinced though... The intercepts will be x = 0 and y = 0 meaning that if he doesn't work out, he won't burn any calories... I'm not convinced because the equation we came up with implies that he is biking and inline skating at the same time.

OpenStudy (anonymous):

hmmm... give me a minute. I won't leave you until we figure this out. Promise :)

OpenStudy (anonymous):

ugh! i hate algebra!! -.-

OpenStudy (anonymous):

I usually don't have a problem with it... Maybe it's just because it's been a really long day. We will get this :)

OpenStudy (anonymous):

Maybe if we work backwards... If he inline skated for 30 minutes, he burned 30 * 7.5 calories, right?

OpenStudy (anonymous):

making it 225, then 420 - 225 = 195 calories left meaning he biked 19.5 minutes (195/10)

OpenStudy (anonymous):

Yeah true! but idk somehow it just still doesnt seem quite right

OpenStudy (anonymous):

ohhh i think i got something.

OpenStudy (anonymous):

If we do an input/output table, then we get: